OpenStudy (anonymous):

anyone know what 33% of 65 1/8 is? I will medal and fan :)

OpenStudy (texaschic101):

0.33(65.125) = 21.49125 or 0.33 * 521/8 = 171.93/8 = 21.49125

OpenStudy (anonymous):

thanks :)